6. Independent events A and B are such that P(A) = a, P(B) = 2a and P(A U B) = 5/8. Find P(A) and P(B).
Answer #1

If two events A and 2B are independent, P(A) x P(B) = P(A \cap B)

P(A \cap B) = P(A) + P(B) - P(A U B)

P(A) x P(B) = P(A) + P(B) - P(A U B)

a x 2a = a + 2a - 5/8

2a2 - 3a -5/8 = 0

a = 0.2515

P(A) = 0.2515

P(B) = 0.5030